OpenStudy (anonymous):

Solve the equation: x^2-100 = 0 Select one: a. x = 10, -10 b. x = 50, -50 c. x = 0, 10 d. x = 0, -10

OpenStudy (anonymous):

Any thoughts on what to do first?

OpenStudy (anonymous):

When doing algebra and you have an unknown like (x). You have to isolate the unkown on one side of the equation to start.

OpenStudy (anonymous):

So how would we do that?

OpenStudy (anonymous):

by doing x^2?

OpenStudy (anonymous):

We have to isolate any unkown on one side of the equation. X^2 is our unknown. So we need to do some thing with the -100. What would the equation look like if you add 100 to both sides?

OpenStudy (anonymous):

x^2-0=100?

OpenStudy (anonymous):

So we have X^2=100 correct?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Now we are trying to solve for x so how do we \[x ^{2}=x\]

OpenStudy (anonymous):

What would you have to do to x^2 to make it x?

OpenStudy (anonymous):

would the answer be D

OpenStudy (anonymous):

No

OpenStudy (anonymous):

my bad it would be A. cus i combined the like terms and simplified it

OpenStudy (haseeb96):

x^2 -100=0 x^2 =100 now take square root on both sides so it will be x = 10 , -10 is the correct answer u can check it by putting one by one both the term in the given equation in the place of x and if u will get same vale like 0=0 or 1=1 etc then it is correct

OpenStudy (haseeb96):

now it is yur job how to solve it if u have any need of me then tell me

OpenStudy (anonymous):

Great job